# Quantifying Classifier Utility under Local Differential Privacy

### v2

@@ 3.2 Further Steps
4 Formal Framework
This section presents the formal framework for quantifying utility of classifiers under various types of LDP mechanisms.
4.1 Concentration Analysis of LDP
@@
Let $\mathcal{M}(x)$ be an $\varepsilon$-LDP mechanism applied to an input $x\in\mathbb{R}$, and let $F_{\mathcal{M}}$ denote the CDF of the perturbed output $\mathcal{M}(x)$. The concentration property of $\mathcal{M}(x)$ is given by:
@@
This proposition follows directly from the definition of the CDF and provides a probabilistic characterization of the “boundedness” of the LDP mechanism $\mathcal{M}(x)$. As a special case, if $a=x-\theta$ and $b=x+\theta$, it becomes $\Pr[\mathcal{M}(x)\in B_{\theta}(x)]$. For practical mechanisms designed for better utility, $\mathcal{M}(x)$ is typically concentrated around the input $x$. This implies that $F_{\mathcal{M}}(b)-F_{\mathcal{M}}(a)$ captures most of the probability mass when $x\in[a,b]$. We analyze the concentration properties of typical continuous and discrete LDP mechanisms in the following subsections.
4.1.1 Continuous Mechanisms
@@
Noise-adding mechanisms are input-independent, using the same noise distribution for all inputs. Examples include the Laplace and Gaussian mechanisms [17].
@@
where $F_{\mathcal{M}}$ is the CDF of the noise distribution. The last equality holds due to the symmetry of the noise distribution, making the result independent of $x$. For instance, the Laplace mechanism’s concentration analysis is given by:
@@
Piecewise-based mechanisms are input-dependent, meaning the sampling distribution changes for different inputs $x$. Examples include PM [54] and SW [37], which use different $p_{\varepsilon}$ and intervals $l_{x,\varepsilon},r_{x,\varepsilon}]$. New mechanisms can be designed by modifying these parameters.
Concentration Analysis. For a piecewise-based mechanism $\mathcal{M}(x)$, the probability of outputting a perturbed value $\mathcal{M}(x)\in B_{\theta}(x)$ depends on $\theta$. The concentration analysis is given by:
@@
This analysis is $x$-dependent due to the terms $l_{x,\varepsilon}$ and $r_{x,\varepsilon}$. For example, in the PM mechanism with $p_{\varepsilon}=\exp(\varepsilon/2)$, if $B_{\theta}(x)\subseteq[l_{x,\varepsilon},r_{x,\varepsilon}],$ the probability is $2\theta\cdot\exp(\varepsilon/2)$. For $\varepsilon=2$ and $\theta=0.3$, the concentration probability $\mathcal{M}(x)\in B_{\theta}(x)$ is $0.86$, indicating that the PM mechanism is more concentrated than the Laplace mechanism under these parameters.
4.1.2 Discrete Mechanisms
@@
Examples of such mechanisms include the $k$-$R$ mechanism [31], where $p_{\varepsilon}(\tilde{x}_{i},x)$ is uniform except for $\tilde{x}_{i}=x$, and the Exponential mechanism [17], where $p_{\varepsilon}(\tilde{x}_{i},x)$ depends on the distance between $\tilde{x}_{i}$ and $x$.
Concentration Analysis. For discrete mechanisms, the support is a finite set $\mathcal{X}$, and the CDF is defined as: $F_{\mathcal{M}}(\tilde{x})=\sum_{\tilde{x}_{i}\leq\tilde{x}}p_{\varepsilon}(\tilde{x}_{i},x)$. The probability of $\mathcal{M}(x)$ outputting a value in $B_{\theta}(x)$ is:
@@
$p_{\varepsilon}(\tilde{x}_{i},x)=\frac{\exp(\varepsilon)}{|\mathcal{X}|-1+\exp(\varepsilon)}\text{ if }\tilde{x}_{i}=x\text{ o/w }\frac{1}{|\mathcal{X}|-1+\exp(\varepsilon)}.$
If $\mathcal{X}$ discretizes $[0,1]$ evenly, i.e. $\mathcal{X}=\{0,1/|\mathcal{X}|,2/|\mathcal{X}|,\ldots,1\}$, the concentration analysis becomes:
@@
4.1.3 Comparison of Concentration Properties
@@
4.2 Robustness Analysis of Classifier
We analyze the robustness of classifiers under the concept of probabilistic robustness, a probabilistic approximation of deterministic robustness. [^3] If the classifier is white-box, deterministic robustness can be derived analytically or through verification methods. See Appendix C.1 for details. In the most ideal case, the deterministic robustness can be expressed as a closed-form function of arbitrary $x$. We omit such easier cases for generality. Probabilistic robustness suffices for utility analysis under LDP mechanisms, as both the robustness and concentration properties of LDP mechanisms are probabilistic. Compared to deterministic robustness, it is computationally efficient and applicable to any classifier, including black-box classifiers, thus suitable as a general framework.
4.2.1 Probabilistic Robustness
@@
Definition 8 (Probabilistic robustness [52, 64]).
@@
This definition implies that the event $h(\tilde{x})=h(x)$ is almost always true (within a tolerance $\tau$) with probability at least $1-\omega$. Here, $1-\omega$ also corresponds to the confidence level in the hypothesis testing. For minimal $\tau$ and $\omega$, e.g. $\tau=0.01$ and $\omega=0.05$, $h$ is nearly $\theta$-locally robust at $x$. Deterministic robustness is a special case with $\tau=0$ and $\omega=0$.
Probabilistic robustness is typically computed using the Hoeffding bound [52, 64], which provides a probabilistic guarantee for the mean of a random variable. Let $Z$ be the indicator function of $h(\tilde{x})=h(x)$, i.e. $Z=1$ if robust and $Z=0$ otherwise. Then, the empirical mean $\hat{\mu}_{Z}$ (robust frequency) is calculated by sampling $n$ points, and the true robustness probability $\mu_{Z}$ can be derived using the Hoeffding bound.
Theorem 1 (Hoeffding bound [29]).
For any $\tau\geq 0$ and $\omega>0$, the probability $\omega$ that the empirical mean ($\hat{\mu}_{Z}$) is within $\tau$ of the true mean ($\mu_{Z}$) is bounded by:
@@
The number of samples $n$ balances utility guarantees and computational cost. A smaller $\omega$ (more deterministic guarantee) requires more samples $n$. Given $\tau$ and $\omega$, we check if the empirical mean $\hat{\mu}_{Z}$ on $n(\omega,\tau/2)$ samples satisfies $\hat{\mu}_{Z}\in[1-\tau/2,1]$. If this is true, by the Hoeffding bound, $|\mu_{Z}-\hat{\mu}_{Z}|\leq\tau/2$, which implies $\mu_{Z}\in[1-\tau,1]$. Thus, Definition 8 holds, and we can claim that $h$ is robust under $B_{\theta}(x)$ with probability at least $1-\omega$.
@@
4.2.2 Procedure for Finding $\theta$
Figure 3: Robustness radius $\theta$ of a 22D classifier at $x$ and the tested decision boundary by brute force.
@@ 4.2.2 Procedure for Finding $\theta$
Figure 4: Comparison of the robustness probability $\rho(\varepsilon,\theta)$ for different LDP mechanisms with $\varepsilon=2,4,6$. Details of the mechanisms and discussions on the curves of $\rho(\varepsilon,\theta)$ are provided in Appendix C.2.
@@
In the above example, any perturbation within the robustness region $B_{\theta}(x)$ ensures that the classifier’s output remains unchanged, thereby preserving utility. When the perturbation is performed by an LDP mechanism, the perturbed value $\tilde{x}$ is drawn from $B_{\theta}(x)$ with a specific probability. The next subsection establishes the connection between $B_{\theta}(x)$ and this probability to quantify the classifier’s utility under LDP mechanisms.
4.3 Utility Quantification
By combining the concentration analysis of LDP mechanisms and the robustness analysis of classifiers, we derive a utility quantification for a given classifier under an LDP mechanism w.r.t. any privacy parameter $\varepsilon$. This quantification framework provides a systematic evaluation of classifier performance under LDP mechanisms. In contrast, the empirical approach requires re-evaluation for different $\varepsilon$ and fails to establish analytical relationships between utility and $\varepsilon$.
Formally, from Definition 8, the classifier is robust under $B_{\theta}(x)$ within tolerance $\tau$ with probability $(1-\omega)$. Assuming the remaining $\omega$ fraction (and similarly, the $\tau$ fraction) of inputs are entirely non-robust, a conservative probabilistic estimate of the exact robustness (without tolerance) is $(1-\omega)(1-\tau)$. Meanwhile, the LDP mechanism $\mathcal{M}$ perturbs $x$ into $B_{\theta}(x)$ with probability $F_{\mathcal{M}}(x+\theta)-F_{\mathcal{M}}(x-\theta)$. Thus, the utility guarantee of the classifier under the LDP mechanism is quantified by the product of these two probabilities. Specifically, given a $d$-dimensional classifier $h$ and raw data $x$, denote $\mathcal{M}^{d}(x)$ as the perturbed $d$-dimensional input by $d$ independent $\mathcal{M}$. [^5] For simplicity, we assume $\mathcal{M}$ is applied to all dimensions of $h$. In practice, it may only be applied to a subset of sensitive features. We claim: With probability at least $\rho(\varepsilon,\theta)$, the classifier $h$ preserves the correct classification result under the input perturbed by $\mathcal{M}^{d}$, where
$\begin{split}\rho(\varepsilon,\theta)&=\Pr[\mathcal{M}^{d}(x)\in B_{\theta}(x)]\cdot\Pr_{\tilde{x}\sim B_{\theta}(x)}[h(\tilde{x})=h(x)]\\ &\geq\left(F_{\mathcal{M}}(x+\theta)-F_{\mathcal{M}}(x-\theta)\right)^{d}\cdot(1-\omega)(1-\tau),\end{split}$
@@ 4.3 Utility Quantification
Furthermore, given the distribution of $x$ as $P_{x}$, we can extend the utility quantification to provide average-case (expected) and worst-case guarantees:
@@
Such guarantees provide input-distribution-aware utility quantification for the classifier. For example, the average-case utility guarantee can be stated as follows: On average, with probability $\rho_{\mathrm{avg}}(\varepsilon)$, the classifier $h$ preserves the correct classification result for an input $x\sim P_{x}$ perturbed by $\mathcal{M}^{d}$.
@@
Takeaway results. Although no mechanism is universally optimal for all classifiers (with a given robustness radius $\theta$) under all $\varepsilon$, heuristic guidelines can assist in mechanism selection. From the above example and Figure 4, we observe that if the privacy parameter $\varepsilon$ is not too small ($\geq 2$), and the classifier’s robustness radius $\theta$ is not too small ($\geq 0.1$), then the PM mechanism is generally the best choice. Experiments in Section 7 further support this conclusion for higher-dimensional classifiers.
Application 2: Select $\varepsilon$ for a utility requirement. With a chosen mechanism, the utility quantification can guide the selection of the privacy parameter $\varepsilon$ to meet a desired utility guarantee. For example, if the classifier’s robustness radius $\theta$ is $0.3$, and we require the classifier to preserve the correct classification result with probability $\rho(\varepsilon,\theta)\geq 0.8$ under the PM mechanism, we can set $\varepsilon=1.8$.
@@
5 Refinement Techniques
This section introduces two techniques to refine the utility quantification presented in Section 4.3. The key observation is that the utility guarantee $\rho(\varepsilon,\theta)$ is a conservative lower bound. It is an at-least guarantee based on the robustness radius under pure $\varepsilon$-LDP. To refine it, we propose the following techniques.
@@ 5 Refinement Techniques
  • Privacy relaxation. We relax the pure $\varepsilon$-LDP constraint to $(\varepsilon,\delta)$-PAC LDP, which tolerates a small probability of failure. Under this relaxation, we introduce a $\delta$-probabilistic indicator for applying the $\varepsilon$-LDP mechanism and propose an extended Gaussian mechanism tailored for PAC LDP.
5.1 Robustness Hyperrectangle

### v3

@@ 3.2. Further Steps
4. Formal Framework
This section presents the formal framework for quantifying classifier utility under input perturbations from various types of LDP mechanisms.
4.1. Concentration Analysis of LDP
@@
Let $\mathcal{M}(x)$ be an $\varepsilon$-LDP mechanism applied to data $x\in\mathbb{R}$, and let $F_{\mathcal{M}}$ denote the CDF of the perturbed data $\mathcal{M}(x)$. The concentration property of $\mathcal{M}(x)$ is given by:
@@
This proposition follows directly from the definition of the CDF and provides a probabilistic characterization of the “boundedness” of the LDP mechanism $\mathcal{M}(x)$. As a special case, if $a=x-\theta$ and $b=x+\theta$, it becomes $\Pr[\mathcal{M}(x)\in B_{\theta}(x)]$. For practical mechanisms designed for better utility, $\mathcal{M}(x)$ is typically concentrated around $x$. This implies that $F_{\mathcal{M}}(b)-F_{\mathcal{M}}(a)$ captures most of the probability mass when $x\in[a,b]$. We analyze the concentration properties of typical continuous and discrete LDP mechanisms in the following subsections.
4.1.1. Continuous Mechanisms
@@
Noise-adding mechanisms are input-independent, using the same noise distribution for all inputs. Examples include the Laplace and Gaussian mechanisms (DBLP:journals/fttcs/DworkR14).
@@
where $F_{\mathcal{M}}$ is the CDF of the noise distribution. The last equality holds because of the symmetry of the noise distribution, making the result independent of $x$. For instance, the Laplace mechanism’s concentration analysis is given by:
@@
Piecewise-based mechanisms are input-dependent, meaning the sampling distribution changes for different inputs $x$. Examples include PM (DBLP:conf/icde/WangXYZHSS019) and SW (DBLP:conf/sigmod/Li0LLS20), which use different $p_{\varepsilon}$ and intervals $l_{x,\varepsilon},r_{x,\varepsilon}]$. New mechanisms can be designed by modifying these parameters.
Concentration Analysis. For a piecewise-based mechanism $\mathcal{M}(x)$, the probability of getting a perturbed value $\mathcal{M}(x)\in B_{\theta}(x)$ depends on $\theta$. The concentration analysis is given by:
@@
This analysis is $x$-dependent due to the terms $l_{x,\varepsilon}$ and $r_{x,\varepsilon}$. For example, in the PM mechanism with $p_{\varepsilon}=\exp(\varepsilon/2)$, if $B_{\theta}(x)\subseteq[l_{x,\varepsilon},r_{x,\varepsilon}],$ the probability is $2\theta\cdot\exp(\varepsilon/2)$. Concretely, for $\varepsilon=2$ and $\theta=0.3$, the concentration probability $\Pr[\mathcal{M}(x)\in B_{\theta}(x)]$ equals $0.86$, indicating that the PM mechanism is more concentrated than the Laplace mechanism under these parameters.
4.1.2. Discrete Mechanisms
@@
Examples of such mechanisms include the $k$-$R$ mechanism (DBLP:conf/nips/KairouzOV14), where $p_{\varepsilon}(\tilde{x}_{i},x)$ is uniform except for $\tilde{x}_{i}=x$, and the Exponential mechanism (DBLP:journals/fttcs/DworkR14), where $p_{\varepsilon}(\tilde{x}_{i},x)$ depends on the distance between $\tilde{x}_{i}$ and $x$.
Concentration Analysis. For discrete mechanisms, the support is a finite set $\mathcal{X}$, and the CDF is defined as: $F_{\mathcal{M}}(\tilde{x})=\sum_{\tilde{x}_{i}\leq\tilde{x}}p_{\varepsilon}(\tilde{x}_{i},x)$. The probability of $\mathcal{M}(x)$ producing a value in $B_{\theta}(x)$ is:
@@
$p_{\varepsilon}(\tilde{x}_{i},x)=\frac{\exp(\varepsilon)}{|\mathcal{X}|-1+\exp(\varepsilon)}\text{ if }\tilde{x}_{i}=x\text{ otherwise }\frac{1}{|\mathcal{X}|-1+\exp(\varepsilon)}.$
Then, if $\mathcal{X}$ discretizes $[0,1]$ evenly, i.e. $\mathcal{X}=\{0,1/|\mathcal{X}|,2/|\mathcal{X}|,\ldots,1\}$, the concentration analysis becomes:
@@
4.1.3. Comparison of Concentration Properties
@@
4.2. Robustness Analysis of Classifier
We analyze the robustness of classifiers under the concept of probabilistic robustness, a probabilistic approximation of deterministic robustness. [^4] If the classifier is white-box, deterministic robustness can be derived analytically or through verification methods. See Appendix C.1 for details. In the most ideal case, the deterministic robustness can be expressed as a closed-form function of arbitrary $x$. We omit such easier cases for generality. Probabilistic robustness suffices for classifier utility analysis under LDP-perturbed inputs, as the concentration properties of LDP mechanisms are also probabilistic. Compared to deterministic robustness, it is computationally efficient and applicable to any classifier, including black-box classifiers, thus suitable as a general framework.
4.2.1. Probabilistic Robustness
@@
Definition 0 (Probabilistic robustness (robust_analysis; DBLP:conf/pkdd/ZhangRF22)).
@@
This definition implies that the event $h(\tilde{x})=h(x)$ is almost always true (within a tolerance $\tau$) with probability at least $1-\omega$. Here, $1-\omega$ also corresponds to the confidence level (type I error) in the hypothesis testing. For minimal $\tau$ and $\omega$, e.g. $\tau=0.01$ and $\omega=0.05$, $h$ is nearly $\theta$-locally robust at $x$. Deterministic robustness is a special case with $\tau=0$ and $\omega=0$.
Probabilistic robustness is typically computed using the Hoeffding bound (robust_analysis; DBLP:conf/pkdd/ZhangRF22), which provides a probabilistic guarantee for the mean of a random variable. Let $Z$ be the indicator function of $h(\tilde{x})=h(x)$, i.e. $Z=1$ if robust and $Z=0$ otherwise. Then, the empirical mean $\hat{\mu}_{Z}$ (robust frequency) is calculated by sampling $n$ points, and the true robustness probability $\mu_{Z}$ can be derived using the Hoeffding bound.
Theorem 6 (Hoeffding bound (Hoeffding_inequality)).
For any $\tau\geq 0$ and $\omega>0$, the probability that the empirical mean ($\hat{\mu}_{Z}$) is within $\tau$ of the true mean ($\mu_{Z}$) is bounded by:
@@
The number of samples $n$ balances utility guarantees and computational cost. A smaller $\omega$ (more deterministic guarantee) requires more samples $n$. Given $\tau$ and $\omega$, we check if the empirical mean $\hat{\mu}_{Z}$ on $n(\omega,\tau/2)$ samples satisfies $\hat{\mu}_{Z}\in[1-\tau/2,1]$. If this is true, by the Hoeffding bound, $|\mu_{Z}-\hat{\mu}_{Z}|\leq\tau/2$ holds, which implies $\mu_{Z}\in[1-\tau,1]$. Thus, Definition 5 holds, and we can claim that $h$ is robust under $B_{\theta}(x)$ with probability at least $1-\tau$, and our confidence level is at least $1-\omega$.
@@
4.2.2. Procedure for Finding $\theta$
Figure 3. Robustness radius $\theta$ of a 22D classifier at $x$ and the tested decision boundary by brute force.
@@ 4.2.2. Procedure for Finding $\theta$
Figure 4. Comparison of the robustness probability $\rho(\varepsilon,\theta)$ for different LDP mechanisms with $\varepsilon=2,4,6$. Details of the mechanisms and discussions on the curves of $\rho(\varepsilon,\theta)$ are provided in Appendix C.3.
@@
In the above example, any perturbation within the robustness region $B_{\theta}(x)$ ensures that the classifier’s output remains unchanged, thereby preserving utility. When the perturbation is performed by an LDP mechanism, the perturbed value $\tilde{x}$ will fall within $B_{\theta}(x)$ with a specific probability. The next subsection establishes the connection between $B_{\theta}(x)$ and this probability to quantify the classifier’s utility under LDP-perturbed inputs.
4.3. Utility Quantification
By combining the concentration analysis of LDP mechanisms and the robustness analysis of classifiers, we derive a utility quantification for a given classifier under inputs perturbed by an LDP mechanism w.r.t. any privacy parameter $\varepsilon$. This quantification framework provides a systematic evaluation of classifier performance under LDP-perturbed inputs. In contrast, the empirical approach requires re-evaluation for different $\varepsilon$ and fails to establish analytical relationships between classifier utility and $\varepsilon$.
Formally, by Definition 5, the classifier is robust under $B_{\theta}(x)$ with tolerance $\tau$ and confidence level $1-\omega$. Therefore, under the same confidence level ($1-\omega$), a probabilistic estimate of the event $h(\tilde{x})=h(x)$, i.e. exact robustness without tolerance, is at least $1-\tau$. Meanwhile, the LDP mechanism $\mathcal{M}$ perturbs $x$ into $B_{\theta}(x)$ with probability $F_{\mathcal{M}}(x+\theta)-F_{\mathcal{M}}(x-\theta)$. Thus, the utility guarantee of the classifier under the LDP mechanism is quantified by the product of these two probabilities. Specifically, given a $d$-dimensional classifier $h$ and raw data $x$, denote $\mathcal{M}^{d}(x)$ as the perturbed $d$-dimensional input by $d$ independent $\mathcal{M}$. [^7] For simplicity, we assume $\mathcal{M}$ is applied to all dimensions of $h$. In practice, it may only be applied to a subset of sensitive features. We claim: Under confidence level at least $1-\omega$, with probability at least $\rho(\varepsilon,\theta)$, the classifier $h$ preserves the correct classification result under the input perturbed by $\mathcal{M}^{d}$, where
$\begin{split}\rho(\varepsilon,\theta)&=\Pr[\mathcal{M}^{d}(x)\in B_{\theta}(x)]\cdot\Pr_{\tilde{x}\sim B_{\theta}(x)}[h(\tilde{x})=h(x)]\\ &\geq\left(F_{\mathcal{M}}(x+\theta)-F_{\mathcal{M}}(x-\theta)\right)^{d}\cdot(1-\tau),\end{split}$
@@ 4.3. Utility Quantification
Furthermore, if the distribution of $x$ as $P_{x}$ is known, we can extend the utility quantification to provide average-case (expected) and worst-case guarantees: [^8] Evaluating average-case and worst-case utility requires knowledge of the input distribution $P_{x}$. In most papers, analyses of worst-case utility for their proposed mechanisms (defined on a specific domain) implicitly assume $P_{x}$ is defined on that domain with non-zero probability. In practice, $P_{x}$ can be estimated from historical data or domain knowledge.
@@
Such guarantees provide input-distribution-aware utility quantification for the classifier. For example, the average-case utility guarantee can be stated as follows: On average, under confidence level at least $1-\omega$, [^9] For clarity, we omit the confidence level in subsequent statements, as it is a predefined constant with $1-\omega\approx 1$. with probability at least $\rho_{\mathrm{avg}}(\varepsilon)$, the classifier $h$ preserves the correct classification result for an input $x\sim P_{x}$ perturbed by $\mathcal{M}^{d}$.
@@
Takeaway results. Although no mechanism is universally optimal for all classifiers (with different robustness radius $\theta$) under all $\varepsilon$, heuristic guidelines can assist in mechanism selection. From the above example and Figure 4, we observe that if the privacy parameter $\varepsilon$ is not too small ($\geq 2$), and the classifier’s robustness radius $\theta$ is not too small ($\geq 0.1$), then the PM mechanism is generally the best choice. Experiments in Section 7 further support this conclusion for higher-dimensional classifiers.
Application 2: Select $\varepsilon$ for a utility requirement. With a chosen mechanism, the utility quantification can guide the selection of the privacy parameter $\varepsilon$ to meet a desired utility guarantee. For example, if the classifier’s robustness radius $\theta$ is $0.3$, and we require the classifier to preserve the correct classification result with probability $\rho(\varepsilon,\theta)\geq 0.8$ under the PM mechanism, we can set $\varepsilon=1.8$ to meet this utility requirement while ensuring a strong privacy guarantee.
@@
5. Refinement Techniques
This section introduces two techniques for refining the utility quantification presented in Section 4.3. The utility guarantee $\rho(\varepsilon,\theta)$ is an at-least guarantee based on the robustness radius under pure $\varepsilon$-LDP. To refine it, we propose the following techniques.
@@ 5. Refinement Techniques
  • Privacy relaxation. We relax the pure $\varepsilon$-LDP constraint to $(\varepsilon,\delta)$-PAC LDP, which tolerates a small probability of failure. Under this relaxation, we introduce a $\delta$-probabilistic indicator for applying the $\varepsilon$-LDP mechanism and propose an extended Gaussian mechanism tailored for PAC LDP. [^10] The primary motivation for adopting PAC LDP is to accommodate the Gaussian mechanism, which is widely used in privacy-preserving machine learning but does not satisfy pure LDP.
5.1. Robustness Hyperrectangle